# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ Abstract.
We obtain the following profinitely rigid examples: the complement of the Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link; specific surgeries on one component of these links; the complement of (full) twist knots $\mathcal{K}_{n}$, Eudave-Muñoz knots $K(3,1,n,0)$, Pretzel knots $P(-3,3,2n+1)$, $5_{2}$ knot; the Berge manifold, and many more.
@@ 1. Introduction
The collection of all finite quotient groups of a finitely generated group $G$, denoted by $\mathcal{C}(G)$, reflects information of the original group $G$. When $G=\pi_{1}M^{3}$ is a 3-manifold group, $\mathcal{C}(G)$ corresponds to the lattice of finite regular coverings of $M$, i.e. the collection of all finite deck transformation groups.
@@
The questions of profinite rigidity in seven of the eight geometries proposed by Thurston have been profoundly understood. For instance, $Sol$-manifolds are profinitely “almost” rigid [26], but not profinitely rigid by [44, 20]; in fact, the profinite rigidity in $Sol$-manifolds can be reduced to a pure number-theoretic problem. [47], together with [57, Corollary 8.3] for the bounded case, provides a complete profinite classification for orientable Seifert fibered spaces. Thus, the most challenging part in profinite rigidity of 3-manifolds is the hyperbolic case.
Liu [35] first showed that the profinite completion of the fundamental group of a finite-volume hyperbolic 3-manifold (including both closed and cusped hyperbolic manifolds) determines its homeomorphism type up to finitely many possibilities. Yet, whether profinite rigidity holds generally in the hyperbolic case is still unknown. There are a few examples of profinitely rigid closed hyperbolic 3-manifolds, for instance the Weeks manifold [8], the $\mathrm{Vol(3)}$ manifold [10], and 00-surgery of the knots ${6_{2}}$ and ${6_{3}}$ in the Rolfsen table by [15]. Some families of fibered hyperbolic 3-manifolds with cusps are also known to be profinitely rigid among all compact 3-manifolds. These examples include hyperbolic once-punctured torus bundles over $S^{1}$ by [11], including the figure-eight knot complement [9]; and hyperbolic four-punctured sphere bundles over $S^{1}$ by [17], including the “magic” manifold. The proof of profinite rigidity for these families are actually based on the congruence omnipotence property of the corresponding mapping class groups.
@@
Suppose $M$ and $N$ are orientable cusped finite-volume hyperbolic 3-manifolds, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then there exists a homeomorphism $\Psi:\partial M\to\partial N$ such that for any boundary slopes $(c_{i})$ on $\partial M$ (allowing empty slopes), there is an isomorphism $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\Psi(c_{i}))}}$ which respects the peripheral structure.
@@
In particular, the boundary homeomorphism $\Psi$ sends exceptional slopes to exceptional slopes.
@@
    The complement of the two-bridge link associated to the rational number $\frac{3}{10}$, denoted by $S^{3}\setminus\mathring{N}(\mathcal{L})$ (Figure 1(c)).
@@
    The complement of the Eudave-Muñoz knots $K(3,1,n,0)$, $n\in\mathbb{Z}\setminus\{0\}$ (Figure 4(a)); in particular, $K(3,1,1,0)$ is the Pretzel knot $P(-2,3,7)$ (Figure 4(b)).
@@
    $5,-1,-2,-4$, and $\frac{5}{2}$-Dehn surgery on one component of the Whitehead link shown in Figure 1(a).
@@
    00-surgery on the $L_{B}^{(1)}$ component of the link $L_{B}$ shown in Figure 5(b).
@@
    00-surgery on the $L_{B}^{(1)}$ component and $r$-surgery on the $L_{B}^{(2)}$ component of the link $L_{B}$ (Figure 5(b)), where $r\in\mathbb{Q}\setminus\{0,4\}$.
@@
Figure 6. The tangle $\mathcal{Q}$ in $S^{2}\times I$, [24, Figure 22.14]
@@
  1. (1)
    The twist knot $\mathcal{K}_{-1}$ is in fact the figure-eight knot, and the profinite rigidity of the figure-eight knot complement was originally proven by Bridson and Reid [9].
  2. (2)
    The $L_{B}^{(2)}$ and $L_{B}^{(3)}$ components of the link $L_{B}$ can be viewed as a $4$-braid in the complement of the $L_{B}^{(1)}$ component, which is homeomorphic to $D^{2}\times S^{1}$. Thus, the 00-surgery on the $L_{B}^{(1)}$ component is in fact a four-punctured sphere bundle over $S^{1}$. The profinite rigidity of this manifold then follows from a general result recently proven by Cheetham-West [17].
  3. (3)
    The integral Dehn surgeries $\mathcal{W}(-1)$, $\mathcal{W}(-2)$, $\mathcal{W}(-4)$, $\mathcal{W}(5)$ are in fact once-punctured torus bundle over circle according to [29, Proposition 3], and their profinite rigidity follows from a general result by Bridson, Reid and Wilton [11].
However, our method of proof is different from [9, 17, 11], which may provide some new insights.
Remark 1.7.
It is conjectured that the profinite completion of fundamental group of a finite-volume hyperbolic 3-manifold determines its hyperbolic volume. The examples listed in Theorem B include some cusped hyperbolic 3-manifolds with small volumes. For instance, the Whitehead link complement and the Whitehead sister link complement are exactly the minimal volume orientable hyperbolic 3-manifolds with two cusps [1], and $\mathcal{W}(\frac{5}{2})$ is one of the second smallest volume orientable hyperbolic 3-manifolds with one cusp [21].
Knot or link complements provide easily accessible examples for 3-manifolds, and the profinite properties of knot complements have long been a popular topic. Wilkes [50] showed that the fundamental group of the complement of a graph knot is profinitely rigid among all fundamental groups of compact, orientable 3-manifolds, which implies the profinite rigidity of prime graph knot complements among all knot complements in $S^{3}$. Ueki [46] showed that that if two knot complements in $S^{3}$ are profinitely isomorphic, then they have the same Alexander polynomial. Cheetham-West [16] introduced a criteria for the profinite rigidity of a hyperbolic knot complement among all knot complements in $S^{3}$, through examining the characterising property and the profinite rigidity of its 00-surgery.
In fact, profinite properties of knot complements are more computable, providing examples that lucidly illustrates our techniques. As an application of Theorem A, we are able to detect hyperbolic knot complements and hyperbolic-type satellite knot complements through the finite quotient groups, and further to match up the Dehn surgeries of a profinitely isomorphic pair of such knots.
@@
Based on Theorem C, we can improve the conclusions by Cheetham-West [16] in determining the profinite rigidity of knot complements through their characterising slopes.
@@
Let $K\subseteq S^{3}$ be a knot. A slope $\alpha\in\mathbb{Q}$ on $\partial N(K)$ is an unoriented characterising slope for $K$, if for any knot $J\subseteq S^{3}$, the Dehn surgery $J(\alpha)$ is homeomorphic to $K(\alpha)$ implies that $J$ is isotopic to $K$ or its mirror image.
There is a slight difference between the definition of an unoriented characterising slope and the usual definition of a characterising slope. Generally, when referring to a characterising slope, it is required that the homeomorphism $J(\alpha)\cong K(\alpha)$ is orientation-preserving, see Definition 4.7 for details. However, we remove the orientation-preserving restriction here, and add a “unoriented” prefix for distinction. Regardless of this distinction, if 00 is a characerising slope of $K$, then it is also an unoriented characterising slope of $K$, see Lemma 4.8.
@@
Let $K\subseteq S^{3}$ be either a hyperbolic knot, or a hyperbolic-type satellite knot. Suppose
@@
As an application of Corollary D, we show the profinite rigidity of some knot complements in $S^{3}$, based on the examples of 00-characterising slopes listed in [4, 5].
Theorem E.
@@
    the Pretzel knots $P(-3,3,2n+1)$, where $n\in\mathbb{Z}$ (Figure 7(c));
@@
(c) The Pretzel knot $P(-3,3,2n+1)$
Figure 7. The hyperbolic knots in Theorem E
@@
Figure 8. The satellite knots in Theorem E
@@
Cheetham-West [16, Theorem 1.1] showed that the complement of the $5_{2}$ knot, $15n_{43522}$ knot, and Pretzel knots $P(-3,3,2n+1)$ are profinitely rigid among all knot complements in $S^{3}$. We enlarge this scope to all compact, orientable 3-manifolds. In fact, this improvement can be deduced from Theorem C.
Remark 1.11.
$\mathcal{W}(\frac{5}{2})$, $S^{3}\setminus\mathcal{K}_{n}$ ($n\in\mathbb{Z}\setminus\{0,\pm 1\}$) in Theorem B, and $S^{3}\setminus 5_{2}$, $S^{3}\setminus 15n_{43522}$, $S^{3}\setminus P(-3,3,2n+1)$ in Theorem E provide new examples of profinitely rigid but non-fibered cusped hyperbolic 3-manifolds. This demonstrates a completely different flavor in the proof of profinite rigidity of cusped hyperbolic 3-manifolds comparing to [11, 17, 52].
@@
In Section 4, we apply this technique to knot complements in $S^{3}$ and prove Theorem C as a combination of Theorem 4.1, Theorem 4.3 and Theorem 4.5. This important observation in determining profinite rigidity of knot complements leads us to the proof of Corollary D and Theorem E in § 4.3.
@@ Acknowledgement
The author sincerely thanks Yu Huang for highly inspiring discussions in the surrounding topics. The author would also like to thank his advisor Yi Liu for useful comments.
@@ 2.1. Profinite groups
For any two finitely generated groups $G_{1}$ and $G_{2}$, the theorem of Nikolov-Segal [38] implies that any isomorphism between $\widehat{G_{1}}$ and $\widehat{G_{2}}$ as abstract groups is indeed an isomorphism as profinte groups, i.e. a homeomorphism with respect to the product topology. Thus, we shall not distinguish between an abstract isomorphism and a continuous one in the following context, as we are mainly concerning finitely generated groups.
@@ 2.2. Residual finiteness
A group $G$ is called residually finite if the intersection of all its finite-index subgroups is trivial. It is easy to see that the canonical homomorphism $G\to\widehat{G}$ is injective if and only if $G$ is residually finite.
@@
Suppose $M$ is a compact 3-manifold. Then the fundamental group $\pi_{1}M$ is residually finite.
@@ 2.3. Profinite detection of hyperbolicity
In our context, a finite-volume hyperbolic 3-manifold refers to a compact, orientable 3-manifold with empty or incompressible toral boundary, whose interior admits a complete Riemannian metric of constant sectional curvature $-1$.
@@
Moreover, when $M$ is a cusped finite-volume hyperbolic 3-manifold, [56, Proposition 3.1] showed that the conjugacy classes of peripheral subgroups in $\widehat{\pi_{1}M}$ are exactly the conjugacy classes of the maximal closed subgroups isomorphic to ^$\widehat{\mathbb{Z}}^{2}$. Therefore, the number of cusps is exactly the number of such conjugacy classes in $\widehat{\pi_{1}M}$. We conclude these properties as the following proposition.
@@ 2.4. Peripheral structure
Let $M$ be a compact, orientable 3-manifold with boundary. We use the notation $\partial_{i}M$ to denote a boundary component of $M$. Up to a choice of basepoint, the inclusion induces a homomorphism $\iota_{\ast}:\pi_{1}(\partial_{i}M)\to\pi_{1}(M)$, and $\partial_{i}M$ is incompressible if and only if $\iota_{\ast}$ is injective. The image of $\iota_{\ast}$ is a conjugacy representative of the peripheral subgroup corresponding to $\partial_{i}M$. When $\partial_{i}M$ is incompressible, by a slight abuse of notation, we simply denote $\pi_{1}\partial_{i}M$ as a peripheral subgroup in $\pi_{1}M$.
@@
Suppose $M$ is a compact orientable 3-manifold, and $\partial_{i}M\cong T^{2}$ is an incompressible toral boundary component. Then there is a canonical isomorphism $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M$.
@@
According to [27] any abelian subgroup of $\pi_{1}M$ is separable. As a result, $\pi_{1}M$ induces the full profinite topology on $\pi_{1}\partial_{i}M\cong\mathbb{Z}^{2}$. The standard theories of profinite groups [41, Lemma 3.2.6] implies that $\widehat{\pi_{1}\partial_{i}M}\to\widehat{\pi_{1}M}$ is injective. Note that the image of this homomorphism is exactly $\overline{\pi_{1}\partial_{i}M}$. Thus, $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}$. In addition, $\pi_{1}\partial_{i}M$ is free-abelian with rank $2$. According to [57, Lemma 2.9], there is a canonical isomorphism ^$\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M\cong\widehat{\mathbb{Z}}^{2}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}\partial_{i}M}$. ∎
@@
Analogous to Fact 1.2, the profinite completion of fundamental group together with the peripheral structure encodes the same data as the “peripheral pair quotients” of the fundamental group. As is pointed out in [57, Theorem 2.5], $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ by an isomorphism respecting the peripheral structure if and only if $\mathcal{C}_{\ast}(\pi_{1}(M);\iota_{1_{\ast}}(\pi_{1}(\partial_{1}M)),\cdots,\iota_{n_{\ast}}(\pi_{1}(\partial_{n}M)))=\mathcal{C}_{\ast}(\pi_{1}(N);\rho_{1_{\ast}}(\pi_{1}(\partial_{1}N)),\cdots,\rho_{n_{\ast}}(\pi_{1}(\partial_{n}N)))$.

### v3

@@ Abstract.
We obtain infinitely many profinitely rigid examples, including: the complement of the Whitehead link, Whitehead sister link, $\frac{3}{10}$ two-bridge link; specific surgeries on one component of these links; the complement of (full) twist knots $\mathcal{K}_{n}$, Eudave-Muñoz knots $K(3,1,n,0)$, Pretzel knots P(3,3,2n+1)P(-3,3,2n+1), $5_{2}$ knot; the Berge manifold, and many more.
@@ 1. Introduction
The collection of all finite quotient groups of a finitely generated group $G$, denoted by $\mathcal{C}(G)$, reflects information of the original group $G$. When $G=\pi_{1}M^{3}$ is a 3-manifold group, $\mathcal{C}(G)$ corresponds to the lattice of finite regular coverings of $M$, which is in fact the collection of all finite deck transformation groups.
@@
The questions of profinite rigidity in seven of the eight geometries proposed by Thurston have been profoundly understood. For instance, $Sol$-manifolds are profinitely “almost” rigid [27], but not profinitely rigid by [44, 21]; in fact, the profinite rigidity in $Sol$-manifolds can be reduced to a pure number-theoretic problem. [47], together with [56, Corollary 8.3] for the bounded case, provides a complete profinite classification for orientable Seifert fibered spaces. Thus, the most challenging part in profinite rigidity of 3-manifolds is the hyperbolic case.
Liu [35] first showed that the profinite completion of the fundamental group of a finite-volume hyperbolic 3-manifold (including both closed and cusped hyperbolic manifolds) determines its homeomorphism type up to finitely many possibilities. Yet, whether profinite rigidity holds generally in the hyperbolic case is still unknown. There are a few examples of profinitely rigid closed hyperbolic 3-manifolds, for instance the Weeks manifold [9], the $\mathrm{Vol(3)}$ manifold [11], and 0-surgery of the knots $6_{2}$} and $6_{3}$} in the Rolfsen table by [16]. The results for these manifolds are even stronger. In fact, the fundamental groups of these manifolds are profinitely rigid among all finitely generated residually finite groups. As for cusped hyperbolic 3-manifolds, the figure-eight knot complement is proven to be profinitely rigid among all compact 3-manifolds by Bridson-Reid [10].
1.1. Profinite rigidity from Dehn filling
@@
Suppose $M$ and $N$ are orientable cusped finite-volume hyperbolic 3-manifolds, and $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, there exists a homeomorphism $\Psi:\partial M\to\partial N$ such that for any boundary slopes $(c_{i})$ on $\partial M$ (allowing empty slopes), $\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\Psi(c_{i}))}}$; moreover, this isomorphism respects the peripheral structure (see Definition 2.5).
Combining Theorem A with the aforementioned results on profinite detection of geometrization, we obtain the following corollary.
@@
In the context of Theorem A, the boundary homeomorphism $\Psi$ sends exceptional slopes bijectively to exceptional slopes.
@@
    The complement of the two-bridge link with Schubert normal form $\frac{3}{10}$, denoted by $S^{3}\setminus\mathring{N}(\mathcal{L})$ (Figure 1(c)).
@@
    The complement of the Eudave-Muñoz knots $K(3,1,n,0)$, $n\in\mathbb{Z}\setminus\{0\}$ (Figure 4(a)); in particular, $K(3,1,1,0)$ is the Pretzel knot P(2,3,7)P(-2,3,7) (Figure 4(b)).
@@
    5,1,2,45,-1,-2,-4, and $\frac{5}{2}$-Dehn surgery on one component of the Whitehead link shown in Figure 1(a).
@@
    0-surgery on the $L_{B}^{(1)}$ component of the link $L_{B}$ shown in Figure 5(b).
@@
    0-surgery on the $L_{B}^{(1)}$ component and $r$-surgery on the $L_{B}^{(2)}$ component of the link $L_{B}$ (Figure 5(b)), where $r\in\mathbb{Q}\setminus\{0,4\}$.
@@
Figure 6. The tangle $\mathcal{Q}$ in $S^{2}\times I$, [25, Figure 22.14]
@@
It is conjectured that the profinite completion of the fundamental group of a finite-volume hyperbolic 3-manifold determines its hyperbolic volume. The examples listed in Theorem B include some cusped hyperbolic 3-manifolds with small volumes. For instance, the figure-eight knot complement (which is the knot $\mathcal{K}_{-1}$ in (4) and also the 1-1-surgery on the Whitehead link in (8)) and the figure-eight sibling manifold (which is the $5$-surgery on the Whitehead link in (8)) are exactly the two minimal volume one-cusped orientable hyperbolic 3-manifold [14]. The Whitehead link complement in (1) and the Whitehead sister link complement in (2) are exactly the minimal volume two-cusped orientable hyperbolic 3-manifolds [1]. The $\frac{5}{2}$-surgery on the Whitehead link is one of the second smallest volume one-cusped orientable hyperbolic 3-manifold, and the 2-2-surgery on the Whitehead link is one of the third smallest volume one-cusped orientable hyperbolic 3-manifold [22].
1.2. Application to knot complements
Knot or link complements provide easily accessible examples for 3-manifolds, and the profinite properties of knot complements have long been a popular topic. Boileau-Friedl [7, Theorem 1.5] showed that the complement of any torus knot is profinitely rigid among all knot complements in $S^{3}$. In fact, its fundamental group is profinitely rigid among the fundamental groups of all compact orientable 3-manifolds by [56, Corollary 8.3]. Similar results were proved for graph knots by Wilkes [50] that the fundamental group of the complement of any graph knot is profinitely rigid among the fundamental groups of all compact orientable 3-manifolds, which implies the profinite rigidity of prime graph knot complements among all knot complements. Some knot invariants are also proven to be detected by the profinite completion of the knot group, including fiberedness and genus by Boileau-Friedl [7, Theorem 1.2], and the Alexander polynomial by Ueki [46].
Profinite properties of knot complements are more computable, and they provide examples that lucidly illustrates our techniques. As an application of Theorem A, we are able to detect hyperbolic knot complements and hyperbolic-type satellite knot complements through the finite quotient groups, and further to match up the Dehn surgeries of a profinitely isomorphic pair of such knots.
@@ 1.2. Application to knot complements
Cheetham-West [17] found a sufficient condition for a hyperbolic knot complement to be profinitely rigid among all knot complements in $S^{3}$, through examining the characterising property and the profinite rigidity of its 0-surgery. Based on Theorem C, we can improve this criteria.
@@
Let $K\subseteq S^{3}$ be a knot. A slope $\alpha\in\mathbb{Q}$ on $\partial N(K)$ is an unoriented characterising slope for $K$, if for any knot $J\subseteq S^{3}$, the Dehn surgery $J(\alpha)$ being homeomorphic to $K(\alpha)$ implies that $J$ is isotopic to $K$ or its mirror image.
There is a slight difference between the definition of an unoriented characterising slope and the usual definition of a characterising slope. Generally, when referring to a characterising slope, it is required that the homeomorphism $J(\alpha)\cong K(\alpha)$ is orientation-preserving, see Definition 4.7 for details. However, we remove the orientation-preserving restriction here, and add a “unoriented” prefix for distinction. Regardless of this distinction, if 0 is a characerising slope of $K$, then it is also an unoriented characterising slope of $K$, see Lemma 4.8.
@@
As an application of Corollary D, we show the profinite rigidity of some knot complements in $S^{3}$, based on the examples listed in [4, 5] where 0 is a characterising slope.
Corollary E.
@@
    the Pretzel knots P(3,3,2n+1)P(-3,3,2n+1), where $n\in\mathbb{Z}$ (Figure 7(c));
@@
(c) The Pretzel knot P(3,3,2n+1)P(-3,3,2n+1)
Figure 7. The hyperbolic knots in Corollary E
@@
Figure 8. The satellite knots in Corollary E
@@
Cheetham-West [17, Theorem 1.1] showed that the complement of the $5_{2}$ knot, $15n_{43522}$ knot, and Pretzel knots P(3,3,2n+1)P(-3,3,2n+1) are profinitely rigid among all knot complements in $S^{3}$. We enlarge this scope to all compact, orientable 3-manifolds. In fact, this improvement can be deduced from Theorem C.
@@
In Section 4, we apply this technique to knot complements in $S^{3}$ and prove Theorem C as a combination of Theorem 4.1, Theorem 4.3 and Theorem 4.5. This important observation leads us to the proof of Corollary D and Corollary E in § 4.3.
@@ Acknowledgement
The author sincerely thanks Yu Huang for highly inspiring discussions in the surrounding topics, and also his advisor Yi Liu for useful comments. He would also like to thank the referee for careful proofreading and valuable suggestions.
@@ 2.1. Profinite groups
$G_{1}$$G_{2}$$^{\widehat{G_{1}}}$$^{\widehat{G_{2}}}$$\scriptstyle{f}$$\scriptstyle{\widehat{f}}$
For any two finitely generated groups $G_{1}$ and $G_{2}$, the theorem of Nikolov-Segal [38] implies that any homomorphism from $\widehat{G_{1}}$ to $\widehat{G_{2}}$ as abstract groups is continuous. Thus, when we say that the profinite completions of two finitely generated groups are isomorphic, in addition to a group-theoretical isomorphism, the isomorphism is automatically a homeomorphism with respect to their topologies defined in the beginning of this subsection.
Lemma 2.1 ([39, Proposition 3.2]).
Let $G_{1}$ and $G_{2}$ be finitely generated groups such that $\widehat{G_{1}}\cong\widehat{G_{2}}$. Then, their abelianizations $G_{1}^{\mathrm{ab}}$ and $G_{2}^{\mathrm{ab}}$ are isomorphic.
2.2. Powers in a profinite group
[…]
Suppose $\Pi^{\prime}\subseteq\Pi$ is a closed subgroup such that $x\in\Pi^{\prime}$. Then, $\Pi^{\prime}$ is also a profinite group by [41, Proposition 2.2.1], and $\varphi_{x}^{\Pi}$ is the composition of $\varphi_{x}^{\Pi^{\prime}}$ with the inclusion map $\Pi^{\prime}\hookrightarrow\Pi$. Thus, for any $\lambda\in\widehat{\mathbb{Z}}$, the $\lambda$-power of $x$ in $\Pi^{\prime}$ is the same as the $\lambda$-power of $x$ in $\Pi$, and we are justified to use the notation $x^{\lambda}$ without specifying the ambient subgroup. We refer the readers to [41, Chapter 4] for more details.
@@ 2.3. Residual finiteness
An abstract group $G$ is called residually finite if the intersection of all its finite-index subgroups is trivial. It is easy to see that the canonical homomorphism $G\to\widehat{G}$ is injective if and only if $G$ is residually finite. Thus, when $G$ is residually finite, we always view $G$ as a subgroup of $\widehat{G}$ via the canonical homomorphism, and any element of $G$ is also regarded as an element of $\widehat{G}$.
@@
The fundamental group of any compact 3-manifold is residually finite.
@@ 2.4. Profinite detection of hyperbolicity
In our context, a finite-volume hyperbolic 3-manifold refers to a compact, orientable 3-manifold with empty or incompressible toral boundary, whose interior admits a complete Riemannian metric of constant sectional curvature 1-1.
@@
Moreover, when $M$ is a cusped finite-volume hyperbolic 3-manifold, [55, Proposition 3.1] showed that the conjugacy classes of peripheral subgroups in $\widehat{\pi_{1}M}$ are exactly the conjugacy classes of the maximal closed subgroups isomorphic to $\widehat{\mathbb{Z}}^{2}$. Therefore, the number of cusps is exactly the number of such conjugacy classes in $\widehat{\pi_{1}M}$. We conclude these properties as the following proposition.
@@ 2.5. Peripheral structure
Let $M$ be a compact, orientable 3-manifold with boundary. We use the notation $\partial_{i}M$ to denote a boundary component of $M$. Up to a choice of basepoint, the inclusion induces a homomorphism $\iota_{\ast}:\pi_{1}(\partial_{i}M)\to\pi_{1}(M)$, and $\partial_{i}M$ is incompressible if and only if $\iota_{\ast}$ is injective. The image of $\iota_{\ast}$ is a conjugacy representative of the peripheral subgroup corresponding to $\partial_{i}M$. When $\partial_{i}M$ is incompressible, by a slight abuse of notation, we sometimes omit the notation $\iota_{\ast}$ and simply denote $\pi_{1}\partial_{i}M$ as a peripheral subgroup in $\pi_{1}M$.
@@
Suppose $M$ is a compact orientable 3-manifold, and $\partial_{i}M\cong T^{2}$ is an incompressible toral boundary component. Then, via the inclusion map, there is an isomorphism $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}^{2}$.
@@
According to [28] any abelian subgroup of $\pi_{1}M$ is separable. As a result, $\pi_{1}M$ induces the full profinite topology on $\pi_{1}\partial_{i}M\cong\mathbb{Z}^{2}$. The standard theories of profinite groups [41, Lemma 3.2.6] implies that $\widehat{\pi_{1}\partial_{i}M}\to\widehat{\pi_{1}M}$ is injective. Note that the image of this homomorphism is exactly $\overline{\pi_{1}\partial_{i}M}$. Thus, $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}^{2}$. ∎
@@
Analogous to Fact 1.2, the profinite completion of fundamental group together with the peripheral structure encodes the same data as the “peripheral pair quotients” of the fundamental group. As is pointed out in [56, Theorem 2.5], $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ by an isomorphism respecting the peripheral structure if and only if $\mathcal{C}(\pi_{1}(M);\iota_{1_{\ast}}(\pi_{1}(\partial_{1}M)),\cdots,\iota_{n_{\ast}}(\pi_{1}(\partial_{n}M)))=\mathcal{C}(\pi_{1}(N);\rho_{1_{\ast}}(\pi_{1}(\partial_{1}N)),\cdots,\rho_{n_{\ast}}(\pi_{1}(\partial_{n}N)))$.